OpenStudy (wes_is_here):

Carrie's Catering is having a special on their box lunches. The minimum purchase is 12 boxes for $120. Each additional box is $8. If x represents the number of boxes over the minimum, which of the following equations represents the cost of box lunches during this special? y = 120x - 8 y = 8x + 120 y = 120 - 8x y = 120x + 8

OpenStudy (swishking19):

It's the second one, since they said the price for each additional box, which was $8.

OpenStudy (swishking19):

So yeah.

OpenStudy (wes_is_here):

can u help with more

OpenStudy (swishking19):

It keeps saying I'm replying even though I'm not lol.

OpenStudy (wes_is_here):

On Fridays at the ballpark, a cup of soda costs $2, and each refill costs $0.70. If x represents the number of refills, which of the following equations represents the amount a person spends on soda? y = 2x + 0.7 y = 2x + 2.7 y = 0.7x + 2.7 y = 0.7x + 2

OpenStudy (wes_is_here):

On Fridays at the ballpark, a cup of soda costs $2, and each refill costs $0.70. If x represents the number of refills, which of the following equations represents the amount a person spends on soda? y = 2x + 0.7 y = 2x + 2.7 y = 0.7x + 2.7 y = 0.7x + 2

OpenStudy (wes_is_here):

?????????